|x - 2| + 10 = 12
thanks for your help

Answer by ankor@dixie-net.com(12696)

|x - 2| + 10 = 12
subtract 10 from both sides
|x - 2| = 12 - 10
:
|x - 2| = 2
remove the absolutes: two solutions:
x - 2 = 2
x = 2 + 2
x = 4
and
x - 2 = -2
x = -2 + 2
x = 0
:
:
Check both solutions in the original equation